OpenStudy (anonymous):

Let \displaystyle f(x) = 4x\sqrt{2-3x}. Find the equation of line tangent to the graph of f at x = -2. not sure how to do this.

zepdrix (zepdrix):

Ok so we want an equation for the line tangent to the graph of f. It will be of the form, \[\Large\bf\sf y_{\tan}\quad=\quad \color{orangered}{m}x+b\] We only have to find a couple of things, the `slope` and the `y-intercept`.

zepdrix (zepdrix):

The derivative evaluated at x=-2 will give us our slope. \[\Large\bf\sf \color{orangered}{m\quad=\quad f'(-2)}\]

zepdrix (zepdrix):

Do you know how to find the slope of f?

OpenStudy (anonymous):

I calculated (or i attempted to calculate) f'(-2) and I got 16(sqrt(2)) but I wasn't sure how to plug that in without b or y since I need y=mx+b.

zepdrix (zepdrix):

If that's correct (I'll check it in a sec :D ), then we plug it in for our m. Then to find b, we need a coordinate pair that we can plug into the tangent line. The function and tangent line will share the coordinates where the tangent line meets the curve. So find a coordinate pair. f(-2) = ?

OpenStudy (anonymous):

I think it's wrong lol I think I forgot to make it -2 when I subbed in.

OpenStudy (anonymous):

wait no. i put in -2

zepdrix (zepdrix):

Hmm I came up with 11sqrt2. Lemme try that again real quick and make sure I didn't mess up somewhere :U

OpenStudy (anonymous):

it's very possible I'm the one who made the mistake I'm not too good with this

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad 4x\sqrt{2-3x}\] \[\Large\bf\sf f'(x)\quad=\quad 4\sqrt{2-3x}+4x\frac{-3}{2\sqrt{2-3x}}\] Evaluating at -2,\[\Large\bf\sf f'(-2)\quad=\quad 4\sqrt{8}+\frac{24}{2\sqrt{8}}\]That look correct so far? :o

OpenStudy (anonymous):

can you tell me how you got the 3 in the numerator? because I had 1/(2sqrt(2-3x))

OpenStudy (anonymous):

for square roots isn't it 1/2sqrtx?

zepdrix (zepdrix):

This is a good derivative to memorize:\[\Large\bf\sf \frac{d}{dx}\sqrt{x}\quad=\quad \frac{1}{2\sqrt x}\] Yes very good. Looks like you remembered that one :) Just have to chain rule after that, multiplying the derivative of the inner function.

OpenStudy (anonymous):

what i did was d/dx (4x)(sqrt2-3x)+(4x)d/dx(sqrt2-3x) and so I assumed that the derivative of sqrt of 2-3x was just (1/(2)(sqrt(2-3x)) I'll have to look back at my notes about the chain rule because I must have forgot

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}\sqrt{2-3x}\quad=\quad \frac{1}{2\sqrt{2-3x}}\cdot\frac{d}{dx}(2-3x)\]Yah we can't forget that chain rule!! :D

OpenStudy (anonymous):

so you put the 2-3x in the numerator too?

OpenStudy (anonymous):

or. derivative of 2-3x

OpenStudy (anonymous):

that's what i meant. derivative.

zepdrix (zepdrix):

Yes, the derivative, to save space :)

OpenStudy (anonymous):

okay that makes sense, i see where the 3 comes in haha.

zepdrix (zepdrix):

Then to find b, you need a coordinate pair.\[\Large\bf\sf f(-2)\quad=\quad 4(-2)\sqrt{2-3(-2)}\quad=\quad ?\]

OpenStudy (anonymous):

so would that be \[16\sqrt{2}\] ?

zepdrix (zepdrix):

Mmm yes! But with a negative in front right?

OpenStudy (anonymous):

yes sorry! i have it in my notes forgot to add that lol

zepdrix (zepdrix):

\[\Large\bf\sf y_{\tan}\quad=\quad 11\sqrt2 x+b\]So to find the y-intercept you plug in the coordinate we found. ( -2, -16sqrt2 )

OpenStudy (anonymous):

so i would put \[16\sqrt{2}\] for b?

OpenStudy (anonymous):

-16*

zepdrix (zepdrix):

Hmm I think it's \(\Large\bf\sf b=6\sqrt2\) Might wanna check your work again :O

OpenStudy (anonymous):

i was just asking what to plug in. So since ( -2, -16sqrt2 ) are the coordinates, would i plug in -2 for x and then 16sqrt2 for y and solve for b?

zepdrix (zepdrix):

Oh yes +_+

OpenStudy (anonymous):

okay i see how you got b now and i also got \[6\sqrt{2}\]

OpenStudy (anonymous):

so the equation would be \[y=11\sqrt{2} +6\sqrt{2}\]?

zepdrix (zepdrix):

With an x in there somewhere, yes? :)

OpenStudy (anonymous):

yes yes i keep forgetting stuff haha. so \[y=11\sqrt{2}x+6\sqrt{2}\]

zepdrix (zepdrix):

Mmmm yah looks right Yay! Good job! \c:/

OpenStudy (anonymous):

thank you so much for your help!

zepdrix (zepdrix):

no prob! :o Oh also I noticed you've only been here a lil while, \(\Large\bf \color{#8C3D62 }{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

thanks :) It's one of the best things I've discovered on the internet!